Determine the solubility of calcium hydroxide in molL-1 and gL-1 (0.01-0.10 molL-1 is


classified as sparingly soluble).

1
Expert's answer
2022-01-25T05:12:02-0500

Ca ( OH )2 = Ca2+ + 2 OH- .


Given Solubility of Ca ( OH )2 = 0.01 mol/ l .


Molar mass of Ca( OH )2 = 40 + 34 = 74 .


Solubility of Ca ( OH )2 = ( 0.01 × 74 ) g / L .

= 0.74 g / L .


Solubility prodict = 4 ( 0.01 )3 . = 4 × 10-6 .